Consider the initial value problem (t-2) y + cot(t) y +ty=e, y( 3 ) = 41/3, ( 3 ) =- T/ 4. Without solving the equation,
Answer #1

# Re-write . D.&. as BCH) = Cost/ (-2) Y + DC+) y + act)y = 70+) ; OCH = t/C&-2) rH= eti(8-2) YCto) = 9. Yo = for Y (3)=437

Reference :

Theorem: Existence and Uniqueness Let p(t), q(t), and g(t) be continuous on [a,b], then the differential equation y + p(t)y